OpenStudy (anonymous):

Algebra Home Work! Will give medal to best answer! 24.) The volume, V, of a gas varies inversely as the pressure, P. If the volume of a gas is 200 L under a pressure of 32 atm, what will be its volume under a pressure of 40 atm? A. 120 L B. 250 L C. 160 L D. 240 L 26. Which of the following situations describes a nonlinear relationship? A. The number of students enrolled in a local college increases by 5% each semester. B. For every question missed on an exam, a student's score decreases by six points. C. The number of graduates with a mathematics degree increases by 20 graduate

OpenStudy (anonymous):

V = k / P, where k is a constant. So plug in to find k. 32 = k / 200 K = 6400 So then: V = 6400 / P V = 6400 / 40 V = 160

OpenStudy (anonymous):

26) A is a nonlinear relationship because it increases by 5% each semester. Think about it like this: if you have 100 students at first, and it increases by 5%, you then have 105 students. But then, 5% of 105 is not five. It is 5.25. Thus, it does not increase linearly. B decreases by exactly 6 pts every time, and C increases by exactly 20 each time. There is no change, so those two are linear.

OpenStudy (anonymous):

Thank You SOOOOOOOOOOO Much! I mean it ! ! !